Colloquium

A Mathematics Inquiry Learning at Kutztown Undergraduate Research Presentation (MILK-URP) Lecture

11:00 a.m. Friday, April 22, 2011

Lytle Hall 214

An Exordium of Self-Intersecting Polygons

Mr. Nicholas Owad

B.S. Mathematics (2010) & B.S. Physics (2010)

Kutztown University of Pennsylvania Graduate

Abstract

In this talk will cover the essential aspects of Twisting Theory, a mechanism that was developed by the author to generate self-intersecting polygons through sequences of twists.

After exploring the basic definitions and operations of the theory, stars and bowties will be presented in depth as examples of the connections between twists and self-intersections. We will state and prove some major theorems, and conclude with open questions and conjectures.

11:00 a.m. refreshments served

11:10 a.m. talk begins

The Mathematics Inquiry Learning at Kutztown (MILK) Lectures are funded by the generous support of the Academy for Inquiry-Based Learning, The Educational Advancement Foundation, and Mr. Harry Lucas, Jr.

Dept. of Mathematics Kutztown Univ. of Pennsylvania (DOMKUP) Dr. Paul S. Ache III, Chair